### v1

# Audit Cards: Contextualizing AI Evaluations

Abstract
AI governance frameworks increasingly rely on audits, yet the results of their underlying evaluations require interpretation and context to be meaningfully informative. Even technically rigorous evaluations can offer little useful insight if reported selectively or obscurely. Current literature focuses primarily on technical best practices, but evaluations are an inherently sociotechnical process, and there is little guidance on reporting procedures and context. Through literature review, stakeholder interviews, and analysis of governance frameworks, we propose “audit cards” to make this context explicit. We identify six key types of contextual features to report and justify in audit cards: auditor identity, evaluation scope, methodology, resource access, process integrity, and review mechanisms. Through analysis of existing evaluation reports, we find significant variation in reporting practices, with most reports omitting crucial contextual information such as auditors’ backgrounds, conflicts of interest, and the level and type of access to models. We also find that most existing regulations and frameworks lack guidance on rigorous reporting. In response to these shortcomings, we argue that audit cards can provide a structured format for reporting key claims alongside their justifications, enhancing transparency, facilitating proper interpretation, and establishing trust in reporting.
1 Introduction
AI governance frameworks [^1] This includes frontier AI safety policies from AI companies (e.g. OpenAI Preparedness Framework), regulations (e.g. EU AI Act), and other norms (e.g. NIST AI Risk Management Framework). See Appendix D for a complete list. are being designed to increasingly rely on audits. [^2] We understand audits as formalized evaluation processes of AI models. Usually, a specific evaluation procedure (e.g. benchmarking, red-teaming, uplift studies) underlies the audit. We refer to these underlying procedures as evaluations. See Appendix A for details. Within these frameworks, audits are meant to (1) identify potential risks, (2) incentivize more responsive development practices, and (3) involve more stakeholders in the system deployment process. However, audits can only fulfill this role effectively if the underlying evaluations are both technically rigorous and effectively integrated into decision-making (Hardy et al., 2024). However, not all evaluations are equally rigorous (e.g., Raji et al., 2022; Birhane et al., 2024; Mökander, 2023; Anderljung et al., 2023; Kolt et al., 2024; Casper et al., 2024; Reuel et al., 2024b; Li & Goel, 2024).
To date, much prior literature has focused on the technical side of audits (see Section 2). However, evaluations are never conducted in a vacuum (Wallach et al., 2025). Their results are intimately shaped, not only by the technical methods employed, but by numerous non-technical details of their context and design (see Section 3). For example, while the research field has long understood the importance of conflict of interest disclosures for academic integrity (e.g., Knerr & D’Amelia, 2020), auditor independence remains a consistent concern in the AI ecosystem (Costanza-Chock et al., 2022; Evans et al., 2023).
Even technically rigorous evaluations can be uninformative or even actively misleading if reported on selectively or obscurely (Ananny & Crawford, 2018). This risk is magnified given the vested interest of AI developers in obtaining favorable audit results. In AI and other fields alike, companies undergoing evaluations have strong incentives to game audits and influence reporting in potentially misleading ways (e.g., Krawiec, 2003; Lu, 2006; Marquis et al., 2016).
[…]
Figure 1: Our audit card template requires reporting on three principles and six features. (Left) The three principles (justification, limitations, and assumptions) further transparency about the process behind an audit. (Right) The six features offer key methodological and contextual information. See Section 3 and Section C.1 for details.
Thus, for AI audits to play a meaningful role in governance, evaluation reporting must be done in a way that minimizes the chance of omitting key contextual information. To address this challenge, we propose “audit cards”. Akin to model and system cards (e.g., Mitchell et al., 2019), audit cards provide a structured format for reporting key background, design choices, and justifications. They are designed to offer a standardized approach to reporting that enhances transparency and facilitates informed interpretations of results. This need not necessarily be reported in a separate audit card document, rather the relevant context can be added to existing reports, such as sections in model and system cards.
Overall, we make four key contributions:
  1. 1.
    Audit cards and survey of evaluation literature: Based on a survey of 25 prior works on AI evaluations (Section 3), we propose a template and checklist (Appendix C) for AI audit cards including three key principles (justification, limitations, and assumptions) and six key types of information (auditor identity, evaluation scope, methodology, resource access, process integrity, and review mechanisms).
  2. 2.
    Survey of frontier AI evaluation reports: We analyze the thoroughness of 20 existing evaluation reports for frontier systems (Section 4).
  3. 3.
    Survey of governance frameworks: We analyze gaps between best reporting practices and 21 existing AI governance frameworks (Section 5).
  4. 4.
    Stakeholder insights: We interview and present perspectives from 10 expert stakeholders on AI evaluation reporting practices (Section 6).
2 Related work
Transparency and reporting: In the past decade, frontier AI research and development have largely shifted from being predominantly academia-driven to predominantly industry-driven (Maslej et al., 2024). This shift has been accompanied by a decrease in transparency around proprietary state-of-the-art systems. As such, increasing transparency and awareness has emerged as a key goal of AI governance (e.g., Felzmann et al., 2020; Haresamudram et al., 2023; Winecoff & Bogen, 2024; Chan et al., 2024; Bommasani et al., 2024; Kolt et al., 2024). One barrier to meaningful transparency is a lack of standards for reporting (Maslej et al., 2024). In response to this challenge, previous works have proposed data, model, system, agent, and usage “cards” (Pushkarna et al., 2022; Mitchell et al., 2019; Gursoy & Kakadiaris, 2022; Casper et al., 2025; Wahle et al., 2023) for documenting key information within the AI ecosystem. In this paper, we build on past work by introducing the notion of an “audit card,” surveying what prior literature suggests they should contain, and analyzing the current state of reporting around audits.
AI audits: Formal evaluations, also known as “audits,” of AI systems have been proposed as a key objective to facilitate transparency and scrutiny (Raji et al., 2022; Anderljung et al., 2023; Mökander, 2023; Li & Goel, 2024; Reuel et al., 2024a). Meanwhile, audits are increasingly incorporated into frameworks for AI governance (e.g., EU AI Act, 2024). However, as we will show in Section 5, current governance frameworks often lack substantial guidance for how to report on audits.
Technically rigorous evaluations: AI systems are audited using a variety of approaches including case studies, benchmarks, red-teaming, and mechanistic analysis (Bengio et al., 2024). However, the science of evaluating AI systems is still nascent (Apollo Research, 2024). Not all evaluations are equally technically rigorous, and their apparent outcomes can be highly sensitive to framing and design (e.g., Schaeffer et al., 2023; Burnell et al., 2023; Khan et al., 2025). Toward improved technical practices, Reuel et al. (2024b) outline a set of 46 best practices for rigorous AI benchmark design. Meanwhile, METR (2024b) work toward a portable standard for framing and conducting capability evaluations. However, unlike prior work, here we focus on rigor in reporting both key technical details and context behind AI audits.
On the inherent sociotechnical nature of evaluations: AI evaluations depend on technical tools to assess system properties and risks. However, they are never conducted in a vacuum; they are always embedded in a broader sociotechnical and political context. Meanwhile, the standards that systems are evaluated against are inherently based on subjective human values. Thus, AI evaluations represent a social science measurement challenge (Wallach et al., 2025). Aside from lacking technical soundness, evaluations can fail to be rigorous and serve the public’s interest for a variety of nontechnical reasons. Prior works have emphasized the role of evaluation integrity and integration in ensuring meaningful oversight (Ojewale et al., 2024; Raji et al., 2022; Sharkey et al., 2024). By surveying prior literature on evaluation procedures (Section 3), analyzing current evaluation reports (Section 4), and analyzing evaluation frameworks (Section 5), we make progress toward a more contextual and critical understanding of AI audits.
3 What information do audit cards need for rigorous reporting?
Methodology: We identify key components of audit cards through an initial literature review of work on AI audits and evaluations, transparency, technical evaluation design, and sociotechnical approaches. [^3] We also refined our audit card template based on structured interviews with expert stakeholders. See Section 6. We selected 25 total papers and manually annotated them, scoring the extent to which the paper recommends that aspect of reporting: is it a central argument of the paper (2), a minor argument of the paper (1), or not mentioned in the paper (0) (see further details in Appendix B). In Table 1, we summarize the perspectives from all 25 papers on the three principles and six features. Next, we expand on these principles and features. See Figure 1 for an overview of an audit card and Section C.2 for an actionable audit card checklist that discusses the relative importance of principles and features for different stakeholders and auditing types.
[…]
Table 1: What contextual details do prior works say are key for audit reporting? We divide reporting into three overarching principles (justifications, assumptions, limitations) and six key features (who, what, how, access, integrity, review). We score each paper’s extent of recommendation of each these as 2 (central argument), 1 (minor argument), and 0 (no mention). Refer to Table 4 for our complete, more granular analysis of papers.
3.1 Overarching principles for transparent reporting
The following three overarching principles behind audit design are key for transparency and methodological clarity. These apply across the entire auditing process and the six specific features. We design audit cards to make it easier to adopt these principles throughout the auditing process. This, in return, enables accurate interpretation, facilitates constructive critique, and builds trust in the evaluation process.
Justifications as explicit arguments to support methodological choices, metric selection, and interpretative frameworks. Explicit reporting on justifications enables process transparency. This justification can be through elements of a rigorous scientific process, such as explaining proxies and experimental coverage (Apollo Research, 2024), or through standardized auditing procedures, such as detailed standards identification (Birhane et al., 2024; Costanza-Chock et al., 2022).
Assumptions to articulate the premises underlying the evaluation design and analysis. Reporting on assumptions clarifies the relationship (or lack thereof) between tests and real-world outcomes, presumptions about what constitutes “good” performance, and details about the threat models considered (Reuel et al., 2024b; Barnett & Thiergart, 2024a; b; Raji et al., 2020).
Limitations to explicitly acknowledge the constraints that affect the validity, reliability, or generalizability of findings. Reporting limitations clarifies constraints in data sampling, potential artifacts in the evaluation process, and boundaries of what auditors can reasonably claim based on the evaluation evidence (Mökander et al., 2024; Barnett & Thiergart, 2024b; Reuel et al., 2024b).
3.2 Specific features for contextual information
Aside from overarching principles, reporting on more specific contextual features of audits helps stakeholders assess trustworthiness. We compile recommendations from past literature into six features for reporting evaluation context in audit cards. For a more granular analysis, see Section C.1, where we further divide these six features into constituent aspects.
  • Who are the auditors: Reporting details of auditor identity facilitates trust and clarifies what potential biases might affect their work. It includes reporting the expertise (relevant domain knowledge, accreditations) (Anderljung et al., 2023; METR, b; Reuel et al., 2024b) and background (track record of conducting evaluations, positionality) of auditors (Costanza-Chock et al., 2022; Gebru et al., 2021). Given worries about the privacy of auditors, information about the background can be anonymized and aggregated.
  • What is evaluated: Reporting on evaluation goals enables clear interpretation of findings. This includes the scope (e.g., model vs. scaffolding) and goal (context of use) (Raji et al., 2020; Birhane et al., 2024; Shevlane et al., 2023), type of evaluation (e.g., capability vs. propensity), as the scope of the evaluation influences what the relevant level of access is (METR, 2024b; Liang et al., 2023), as well as reporting on the evaluation obsolescence criteria. [^4] Examples of obsolescence criteria could include substantial amounts of model fine-tuning, changes in system components, or changes in deployment context. This helps stakeholders understand the shelf life of the findings. It also includes reporting whether the evaluation will be repeated, and if so, at what time or after which conditions are met (Barnett & Thiergart, 2024b; Anderljung et al., 2023).
  • How is it evaluated: Explaining the underlying procedures for the evaluation, such as descriptions of the evaluation setup with justifications of how a capability translates to evaluation goal (Apollo Research, 2024; Raji et al., 2020; Reuel et al., 2024b) and provide interpretation of scores (Dobbe et al., 2021; Selbst et al., 2019). The specific relevant aspects to report for this feature may change depending on the underlying evaluation. We therefore refer to the appropriate literature for the underlying evaluation for a complete list of relevant aspects to report and details on how to report on them, e.g. for benchmarks (Reuel et al., 2024b).
  • Access and resources: Report the constraints under which the evaluation was conducted to allow readers to critically assess the audit’s limitations. This includes, what access auditors have to a system (black/white box, with/without safeguards) and which resources available are available (compute, time, funding) as this can have a large impact on the rigor of evaluations (Barnett & Thiergart, 2024b; Ojewale et al., 2024; Casper et al., 2024).
  • Integrity: Accountability for the auditing process establishes trust in auditors and rules out potential conflicts of interest. This includes reporting transparently about the selection process (how auditors were chosen) and compensation (funding sources, incentive structures). These dynamics can significantly affect the evaluation process and are therefore crucial context (Birhane et al., 2024; Costanza-Chock et al., 2022; Gebru et al., 2021).
  • Review and communication: Reports on if and what independent reviews of audit methods facilitate accountability [^5] We discuss issues of intellectual property and non-disclosure agreements (NDAs) standing in the way of internal review in Section 6 and Section 7. This review can be performed by experts at the organization by cross-checking the final report, insofar as they were not involved in the evaluation beforehand. (Reuel et al., 2024b; Anderljung et al., 2023). This further includes contact information and avenues for corrections, executive summaries (covering the evaluation process and result for non-technical audiences), and clear definitions of key terms (Gebru et al., 2021; Bucknall & Trager, 2023; METR, 2024b; Reuel et al., 2024b).
Some of the above features may be considered confidential in certain contexts, with academic researchers potentially sharing different details than commercial entities (Mitchell et al., 2019). Recommendations about how and with whom audit cards should be shared are beyond the scope of this report. However, redactions offer a simple, well-precedented approach for protecting sensitive information (Boone et al., 2015).
4 What do existing evaluation reports include?
Methodology: To study the thoroughness of existing evaluation reports, we examined 20 reports produced between 2023 and 2025. We selected these 20 due to their focus on frontier systems, ensuring to include reports from both developers and third party organizations. We manually annotated each report using the 3 principles and 6 features from our audit card template (Section 3). For each audit card component, we scored reports on a 0-2 scale for comprehensively reporting that component (2), minimally reporting that component (1), not reporting on this component at all (0). We assessed reporting thoroughness only, not execution quality, making no judgments about how well evaluations were executed. This means, for example, transparently reporting that there was a lack of any review/feedback process and justifying this would receive a full score of 2. We summarize the findings in Table 2.
[…]
Table 2: What contextual details do existing audit reports provide? We score reports as providing comprehensive (2), minimal (1), or no information (0) for each component of an audit card. Table ordered by release date of the report.
While all evaluation reports provide at least a general overview of procedures, the level of detail differs. We find that audit reports consistently offer information on evaluation scope (average score 1.25) and procedures (1.15) used. However, contextual details are much less consistently reported (average 0.86, $\sigma=0.65$) both across organizations and sometimes within the same organization. For example, Google DeepMind released four documents with significantly varying levels of contextual detail (average scores of 1.00, 0.67, 0.56, and 1.44 in chronological order). Only some reports (e.g., Meinke et al. ; Gemma Team et al. (2024); Üstün et al. (2024)) offered detailed information on the scaffolding used to evaluate the system.
Reports inconsistently address assumptions, limitations, and auditor information. While justifications and limitations are more commonly acknowledged (18 of 20) than assumptions (12 of 20), they are typically discussed at a high level rather than relating to specific methodological choices. Ten reports discuss limitations minimally, while eight do so comprehensively. However, there are notable qualitative differences even among more comprehensive discussions. For example, Üstün et al. (2024) discuss limitations in prompting techniques, language transferability, evaluation tool chaining, and reproducibility in significant detail, whereas the OpenAI system card (OpenAI, 2024) only consider limitations of what was being evaluated (e.g. classified information or restricted data) rather than substantial limitations of the procedures. Regarding auditor information, 12 of 20 reports include auditor expertise details, but rarely their background (3 of 20), making an assessment of potential biases in the audit harder.
Reports rarely disclose information about evaluation integrity, resources, review processes, and obsolescence criteria. Only 3 of 20 reports detail organization and auditor selection, conflicts of interest, or contractual arrangements governing the evaluation process. Even in these cases, the report tended to be about the contractual details of human baselines that are part of the evaluation rather than the core auditing team itself. Only 7 of 20 reports specify which resources auditors had access to, including model access level, computational resources, and time constraints. Similarly, only six reported on some form of review and 9 on feedback mechanisms (Meinke et al., ; METR, 2024a; OpenAI, 2024; UK AISI & US AISI, 2024a). Four report whether they underwent any form of peer review or quality assurance process before publication (e.g. Gemini Team et al., 2024; UK AISI & US AISI, 2024b). Five specified an option for public feedback and corrections. Zero reports explicitly present obsolescence criteria.
5 What guidance do current governance frameworks offer?
Methodology: We analyzed key governance frameworks issued by institutions that are considered influential regarding technical evaluations for risks from state-of-the-art models. We selected these frameworks based on jurisdiction, institutional authority, and direct influence on model evaluations (see Appendix D for details). This included official regulations (EU AI Act, US Executive Order), voluntary industry standards (NIST AI Risk Management Framework), and policies from major AI developers. We scored each framework on whether it either explicitly required or recommended reporting of each feature in our audit card template, as shown in Table 3.
[…]
Table 3: What guidance do existing governance frameworks provide for reporting? We give binary scores of 1 (requires or recommends explicit reporting of this feature) or 0 (may mention issue’s importance but not specifically its disclosure or reporting). For company policies, it is scored by whether it is disclosed or not. Citations and details for each framework are in Appendix D.
Existing frameworks offer limited guidance on audit reporting. We find existing frameworks consistently require that models are assessed but do not consistently specify when, by whom, for what, and how. This is somewhat expected, as governance frameworks for emerging technologies tend to crystallize gradually (Linkov et al., 2018). High-level requirements in governance documents are often supplemented with more specific forms of soft law such as codes of practice, guidance, and industry norms. Notably, these frameworks rarely require evaluation reports to disclose matters key to the integrity of the process (only 8 of 21) and the resources and access given to auditors (only 2 of 21).
6 Challenges with auditing according to expert stakeholders
Methodology: To more thoroughly understand audit reporting and perceived gaps between current and best practices, we interviewed 10 experts. We reached out to these experts based on their expertise and familiarity across the evaluation ecosystem. The 10 experts come from a diverse range of private companies, non-profit organizations, and government bodies. Interviewees include: evaluation designers (P1), evaluation developers (P3, P4, P5), evaluation report writers (P1, P3, P5), evaluation-related policy researchers (P6, P7), and those supporting evaluation delivery and standards development (P8, P9). To allow for candid discussions, we committed to working with the level of anonymity that was comfortable for experts. See Section F.1 and Section F.2 for interviewees’ backgrounds and interview questions.
The quality of audits depends greatly on the auditor. Reporting on auditor details may become more critical as the evaluation market grows. Currently, the quality and trustworthiness of evaluators are primarily assessed informally (P1, P6, P7) through impressions of auditors’ work and reputation. Many evaluation reports currently lack detailed information about evaluator selection, training, and reporting methods for various reasons, including time constraints or the belief that technical details about the evaluation sufficiently indicate evaluation quality (P3). However, because third-party auditors’ access to proprietary models is voluntary (P1, P6), regulatory gaps and market failures must be addressed to enable more reporting about auditor selection and engagement terms. Standardized evaluation reporting may become increasingly important as the market expands. Auditing regimes from other industries, such as finance, energy, and medicine, may offer guidance moving forward (P9; see also Anderson-Samways (2024)).
Reporting about the evaluation timeline can help with industry’s standardization efforts. The amount of time spent on designing and executing an evaluation can have a direct effect on the quality and thoroughness of the evaluation. When audits begin and end, and the corresponding versions of model access at each checkpoint, can often be in tension with commercial pressures in companies to develop and deploy quickly (P1, P4, P6). Evaluation resources, including time, are often difficult to measure in a chaotic situation (P3). Standardization and guidance beforehand around auditing timeframes may be greatly appreciated by evaluators (P1, P4, P5, P6). In lieu of laws explicitly specifying timeframes for different stages of evaluation, reporting the time taken by auditors will be crucial to facilitate standardization and coordination.
Evaluation reports are currently used more by auditors than policymakers. Currently, evaluation reports by both internal and external evaluators assume a primarily technical audience. Policymakers are constrained by time, technical understanding, regulatory state of affairs, and often read them at a high level (P6, P9). Thus, it seems that evaluation reports more directly inform the science of evaluations than policy, as they assist with reproducibility and elicitation. In the future, it will likely be useful for government authorities to check compliance with standards. Experts have also said that adapting features to a prioritization for each user group would be valuable (P1, P5, P6, P8). This may be the subject of future research.
7 Discussion and recommendations
There are no established standards around how to report the context of audits. The present lack of standards leads to incompleteness and inconsistency in audit reporting practices. Certain contextual aspects, such as the integrity of auditing processes, are particularly neglected. Lacking transparency around audit methodology and context creates a significant barrier to accountability in the AI ecosystem, where trust in evaluations is essential for informed decision-making by regulators, users, and the public.
Audit cards are a practical solution to reporting that accounts for an audit’s real-world context and constraints. The limitations and inconsistencies in current reporting can limit the ability of evaluations to serve their intended role. As the field grows, having standards for reporting is needed to ensure trust between developers, auditors, regulators, and the public. By offering a checklist for reporting, audit cards are designed to capture essential contextual information and allow for a meaningful interpretation of results.
Audit cards remain one approach among several to increase transparency between developers, users, and other stakeholders. Audit cards are designed to offer a structured checklist applicable to all types of assessment (see Section C.2), but should be viewed as an instrument in a larger AI accountability toolkit. The key context they offer can be helpful to other existing mechanisms, including model and system cards, datasheets, and benchmark details (Mitchell et al., 2019; Gebru et al., 2021; Reuel et al., 2024b). Further research may investigate how these cards can enhance and supplement other transparency methods.
Limitations and concerns:
  • Limited sample size and reporting details. To enable consistent high-level analysis of the current ecosystem, we used simple 2 or 3 point scoring methods. However, this comes at the expense of granular details and nuance, differentiating between reports only to a limited degree. This is particularly pronounced for our analysis of governance frameworks, where we use binary scores. Furthermore, our analysis, while extensive (25 academic papers, 20 evaluation reports, and 21 governance frameworks), is not exhaustive.
  • Trade-offs between reporting quality and audit quality: Implementing thorough reporting through audit cards requires additional time and effort from auditors who often already operate under significant resource constraints. The additional burden of comprehensive reporting may be challenging, particularly for smaller organizations or time-sensitive evaluations.
  • Verification challenges: Even with comprehensive reporting guidelines, there remain challenges in verifying the accuracy of disclosed information. However, fraud is not unique to AI. Other fields handle these challenges through formal scrutiny mechanisms and occasional legal action. Through explicit disclosure, audit cards can be a first step toward a more accountable governance regime.
Key near-term challenges with AI audit reporting include providing rigorous and standardized guidance. In the absence of external, authoritative standards for reporting, stakeholders across the AI ecosystem will struggle to effectively compare and interpret audit results. The EU AI Act (including its Code of Practice) and various contributions from AISIs and NIST, international consortia, labs, and third-party evaluators have offered progress toward a shared understanding of rigor. However, future initiatives may benefit from a higher degree of specificity (e.g., templates) for reporting on the principles and features highlighted in this paper.
Standardized reporting frameworks are key for trustworthy AI audits. The path toward effective AI governance requires both technical innovation and procedural standardization. By establishing shared expectations for auditing transparency through structured reporting mechanisms like audit cards, policymakers can build a more accountable AI governance landscape. Such frameworks ensure evaluations are rigorous, not only in execution, but also in reporting. Ultimately, this enables stakeholders to make informed decisions based on transparent, comparable, and contextually rich information that serves the public interest.
Acknowledgments
This research was supported by the ML Alignment & Theory Scholars (MATS) Program, which provided funding for Leon Staufer and Mick Yang through research stipends. We also thank MATS and our research manager Juan Gil for his organizational assistance and research support.
We express our gratitude to Michael Aird, Lily Stelling, as well as other members of the MATS cohort, participants in the FAR Labs discussion group, and others for their valuable feedback and suggestions throughout the development of this paper.
We are grateful to all participants who contributed through interviews. Their perspectives were instrumental in developing the frameworks presented.
Appendix A Disambiguating auditing and the evaluations process
AI audits, for the purpose of our analysis, refer to formalized evaluation processes of AI models or systems. By formalized, we mean that it is performed systematically and by an organization, i.e., not just being performed ad hoc by individuals. This kind of evaluation is a model audit, as opposed to a governance audit or an application audit (Mökander et al., 2024).
The evaluation is the underlying assessment run on the model during the audit. We account for the fact that evaluations can occur internally or externally, and pre- or post-deployment. Evaluations can assess capabilities, propensities, or risks. Evaluations can be in many forms, such as benchmarks, red-teaming, human uplift studies, and user studies. Our approach considers all these variants, as the contextual features we have identified are relevant to all. It is possible, that in some of these contexts certain features will be more important and should therefore be reported in more detail. We explore the importance of certain features further in Section 7 and Section C.1.
We conceptualize the evaluation process as comprising three main stages: (1) designing the underlying evaluation methodology to achieve its aims (this could be a benchmark, an evaluation suite, or a red-teaming exercise); (2) executing the evaluation and getting results; (3) publishing these results in some form. Our focus is to elaborate on stage (3), that is, what is to be included in these evaluation reports.
When speaking of evaluation reports, we refer to internal or external documents including evaluation results (e.g., a model’s results on a benchmark). Examples of evaluation reports include: (OpenAI, 2024; Meinke et al., ) as well as private evaluation reports produced by third-parties for AI companies and internal reports created by AI companies themselves.
Appendix B Literature review
We reviewed past work on “AI auditing”, “AI evaluations”, “best practices”, “science of evaluations”, “evaluation context” and “reporting”. We then assessed each paper for relevance based off the titles. This only yielded eight papers. We thus expanded to adjacent fields such as transparency and AI auditing, technical evaluation design, and sociotechnical approaches. These fields integrate technical performance metrics with societal assessments and considerations of how technical performance interacts with real-world social, cultural, and institutional factors. We scored recommendations within each paper as central arguments (2), minor arguments (1), and no mention (0). Through four iterations, we refined which aspects were most salient, focusing particularly on aspects that were actionable for auditors. We then categorized and prioritized these aspects based on input from experts. The mapping of aspects to features is shown in Section B.2. The final audit card is depicted in Figure 1 and Appendix C.
B.1 List of papers
  1. 1.
    Towards Publicly Accountable Frontier LLMs (Anderljung et al., 2023)
  2. 2.
    Science of Evals (Apollo Research, 2024)
  3. 3.
    Declare and Justify: Explicit assumptions in AI evaluations are necessary for effective regulation (Barnett & Thiergart, 2024a)
  4. 4.
    What AI evaluations for preventing catastrophic risks can and cannot do (Barnett & Thiergart, 2024b)
  5. 5.
    AI auditing: The Broken Bus on the Road to AI Accountability (Birhane et al., 2024)
  6. 6.
    Structured Access for Third-Party Research (Bucknall & Trager, 2023)
  7. 7.
    Rethink reporting of evaluation results in AI (Burnell et al., 2023)
  8. 8.
    Black-Box Access is Insufficient for Rigorous AI Audits (Casper et al., 2024)
  9. 9.
    A Survey on Evaluation of Large Language Models (Chang et al., 2023)
  10. 10.
    Who Audits the Auditors? Recommendations from a field scan of the algorithmic auditing ecosystem (Costanza-Chock et al., 2022)
  11. 11.
    Hard choices in artificial intelligence (Dobbe et al., 2021)
  12. 12.
    Dimensions of Generative AI Evaluation Design (Dow et al., 2024)
  13. 13.
    Can We Trust AI Benchmarks? (Eriksson et al., 2025)
  14. 14.
    Datasheets for Datasets (Gebru et al., 2021)
  15. 15.
    Devising ML Metrics (Hendrycks & Woodside, 2024)
  16. 16.
    Holistic Evaluation of Language Models (Liang et al., 2023)
  17. 17.
    Task Development Guide (METR, b)
  18. 18.
    Auditing Large Language Models: A Three-Layered Approach (Mökander et al., 2024)
  19. 19.
    Reasons to Doubt the Impact of AI Risk Evaluations (Mukobi, 2024)
  20. 20.
    Towards AI Accountability Infrastructure (Ojewale et al., 2024)
  21. 21.
    Closing the AI Accountability Gap: Defining an End-to-End Framework for Internal Algorithmic Auditing (Raji et al., 2020)
  22. 22.
    BetterBench (Reuel et al., 2024b)
  23. 23.
    Fairness and Abstraction in Sociotechnical Systems (Selbst et al., 2019)
  24. 24.
    Model evaluation for extreme risks (Shevlane et al., 2023)
  25. 25.
    Toward an evaluation science for generative AI systems (Weidinger et al., 2025)
[…]
Table 4: Annotations for all 25 papers (see Section B.1) across the 24 initial contextual aspects (see Section B.2). We score each paper’s discussion of these as 2 (central argument), 1 (minor argument), and 0 (no mention).
B.2 Mapping of initial aspects
[…]
  1. 1.
    Level of access auditors had $\rightarrow$ Access and Resources
  2. 2.
    How did they work with the model developers $\rightarrow$ Integrity
  3. 3.
    Were they trained by the developers to evaluate the model $\rightarrow$ Who are the auditors
  4. 4.
    How much compute given to auditors, which manufacturers/chip model $\rightarrow$ Access and Resources
  5. 5.
    Expertise / background of people involved $\rightarrow$ Who are the auditors
  6. 6.
    What was their initial goal for the evaluation $\rightarrow$ What is evaluated
  7. 7.
    What were their documentation requirements $\rightarrow$ Review and communication
  8. 8.
    What was the contract structure / How were they paid/rewarded $\rightarrow$ Integrity
  9. 9.
    What human input was used / who were they $\rightarrow$ How is it evaluated
  10. 10.
    Peer review process / QA testing $\rightarrow$ Review and communication
  11. 11.
    How does capability or concept translate to benchmark task $\rightarrow$ How is it evaluated
  12. 12.
    Involvement of domain experts during design $\rightarrow$ Who are the auditors
  13. 13.
    Feedback channel (openness + how to contact) $\rightarrow$ Review and communication
  14. 14.
    How should scores be interpreted or used $\rightarrow$ How is it evaluated
  15. 15.
    Assumptions of normative properties are documented $\rightarrow$ Assumptions
  16. 16.
    Non-normative assumptions justified $\rightarrow$ Assumptions
  17. 17.
    Evals as a process should consider eval’s limitations (not that evals are limited) $\rightarrow$ Limitations
  18. 18.
    Interdisciplinary teams / including multiple perspectives $\rightarrow$ Who are the auditors
  19. 19.
    What is the technical goal of the model being evaluated $\rightarrow$ What is evaluated
  20. 20.
    What are the principles behind the development of the model $\rightarrow$ What is evaluated
  21. 21.
    Consider context in which the AI system will be used (e.g. value hierarchies, diverse populations) $\rightarrow$ Assumptions
  22. 22.
    Justification of evaluation methods used / using qualitative evaluations as well $\rightarrow$ Justification
  23. 23.
    Timely, continuous evals $\rightarrow$ What is evaluated
  24. 24.
    Interpretable, easy to read, accessible to a wide audience $\rightarrow$ Review and communication
Appendix C Audit cards
C.1 Details and justifications of features
This section provides an overview of audit card features, including a concise summary of what to report and why this contextual information matters. A detailed list of aspects for each feature can be found in the checklist in Section C.2.
Who are the auditors?
Summary: Documents auditor qualifications through domain expertise, experience, technical familiarity, formal credentials, stakeholder perspectives, and social positionality.
Why it matters: Transparency about qualifications establishes credibility while revealing potential biases. Expertise levels directly affect the ability to assess specific systems, while diverse backgrounds help identify evaluation blind spots. Without this information, readers cannot judge whether auditors possessed the necessary skills to thoroughly evaluate the system or what perspectives might be underrepresented in the analysis.
Relevance across contexts: Basic expertise documentation is essential for all evaluation types. Red-teaming specifically requires specialized adversarial thinking skills and security expertise. In user studies, auditor social positionality becomes particularly important as it may influence how human-AI interactions are interpreted.
What is evaluated?
Summary: Defines evaluation scope, goals, and approach by specifying system version, components, scaffolding used, threat models addressed, and whether capabilities or propensities were evaluated. Includes continuous monitoring plans and obsolescence criteria.
Why it matters: Clear boundaries prevent misinterpretation of results. Understanding threats, contexts, and capability vs. propensity assessment helps readers gauge relevance to their specific concerns. Without specific scope information, readers might incorrectly assume the evaluation covers more than it does, while obsolescence criteria help prevent outdated evaluations from being applied to newer system versions.
Relevance across contexts: Evaluation scope and goal specification are critical across all evaluation types. Benchmarks particularly require clear documentation of obsolescence criteria as AI capabilities rapidly evolve. Safety assessments benefit from explicit threat models tailored to deployment context. Human-centered evaluations need clear articulation of intended user groups and use cases, while deployed systems require specified intervals and conditions for re-evaluation.
How is it evaluated?
Summary: Describes methodology, implementation details, and result interpretation guidance. Explains connections between evaluation methods and capabilities, justifies capability-to-goal translations, and outlines finding limitations. Reporting on this feature can be very extensive, as the specific details for reporting depend on the exact evaluation procedure used. We refer to existing literature on best practices for reporting on evaluation procedures, such as Reuel et al. (2024b); Apollo Research (2024).
Why it matters: Methodology directly impacts result validity and reliability. Understanding how capabilities become measurable criteria and how scores translate to real-world performance prevents misapplication of evaluation outcomes. Without methodological transparency, readers cannot assess whether the evaluation approach actually measures what it claims to measure or understand the confidence level appropriate for various findings.
Relevance across contexts: A description of methodology is essential for all evaluation types. Benchmarks specifically require detailed score interpretation guidance to prevent misapplication of numerical results. Propensity evaluations need robust justification for capability-to-goal translations since they often use proxy measurements. Novel system types need clearer justification for how evaluation approaches measure intended capabilities.
Access and resources
Summary: Details auditor system access level (black/gray/white-box, with/without safeguards), provided documentation, and available computational, time, and financial resources.
Why it matters: Resource constraints and access limitations impact evaluation thoroughness and depth. Limited system access or insufficient resources may prevent auditors from discovering certain vulnerabilities or testing edge cases. Transparency about these constraints helps readers understand which areas received less scrutiny and what types of issues the evaluation might have missed due to practical limitations.
Relevance across contexts: Basic access level information is relevant to all evaluations. Red-teaming findings can differ dramatically between black-box and white-box conditions. Large-scale benchmarks particularly require computational resource documentation.
Integrity
Summary: Addresses auditor selection process, potential “opinion shopping,” conflicts of interest, compensation structures, contract terms, and NDAs. Documents measures ensuring diverse perspectives.
Why it matters: Selection processes and incentive structures reveal potential influences on outcomes. Auditors with financial ties to the evaluated organization may face pressure to produce favorable results, while restrictive NDAs might limit disclosure of critical findings. Transparent disclosure of relationships and contractual limitations builds trust in findings and allows readers to identify potential sources of bias.
Relevance across contexts: Integrity documentation and conflict of interest disclosure is universally important across all evaluation contexts. Selection process transparency becomes critical for high-stakes evaluations where findings may significantly impact deployment decisions. Non-disclosure agreement details matter most when evaluations involve proprietary information that might limit the disclosure of certain findings. Disclosure of prior relationships between auditors and developers matters more for external evaluations to establish credibility and independence.
Review and communication
Summary: Outlines evaluation review process, feedback channels, executive summary quality, key term definitions, and result distribution. Focuses on unambiguous key findings and consistent terminology.
Why it matters: Robust review strengthens findings’ reliability by catching errors and biases. Clear feedback channels enable continuous improvement when new information emerges or vulnerabilities are discovered after publication. Well-defined terminology ensures consistent understanding across different stakeholders, preventing miscommunication about technical concepts and evaluation outcomes that could lead to inappropriate system deployment decisions.
Relevance across contexts: Robust review processes are essential across all evaluation types to ensure reliability and catch potential errors or biases. Interdisciplinary evaluations require clear definitions of domain-specific terminology. Security-sensitive applications need explicit documentation of which findings are shared with which audiences.
C.2 Checklist for audit card
This checklist provides a framework for documenting contextual information about AI system audits. Questions are prioritized with the most important appearing first in each section, though not all questions in a section will be relevant to every evaluation. The checklist serves as a flexible guide rather than a rigid requirement. More comprehensive information generally enhances transparency. Through use of the audit card checklist, it is easier to document evaluation contexts, limitations, and strengths to facilitate informed discussions about AI system capabilities and risks.
  • $\square$
    Who are the auditors
    • $\square$
      Expertise:
      • $\bullet$
        What specific domain expertise (e.g. cybersecurity knowledge) do the auditors possess that is relevant to this evaluation?
      • $\bullet$
        What relevant experience do the auditors have in evaluating similar systems?
      • $\bullet$
        What is the auditors’ familiarity with the specific technology being evaluated?
      • $\bullet$
        Do the auditors hold official accreditations or certifications relevant to this evaluation?
      • $\bullet$
        Have the auditors undergone any training specific to this evaluation (internal or external)?
    • $\square$
      Background:
      • $\bullet$
        What stakeholder perspectives do the auditors represent or have experience with?
      • $\bullet$
        What is the social positionality of the auditors (aggregated demographic information, cultural background)?
  • $\square$
    What is evaluated
    • $\square$
      Scope:
      • $\bullet$
        Which specific version or iteration of the system is being evaluated?
      • $\bullet$
        What specific components of the system are included in the evaluation?
      • $\bullet$
        What aspects of the system were explicitly excluded from the evaluation scope?
      • $\bullet$
        Does the evaluation cover the full system or only specific functionalities?
      • $\bullet$
        What scaffolding (e.g. chain-of-thought prompting, few-shot demonstrations) or additional tools (e.g. web or command line access) were used during the evaluation?
    • $\square$
      Goal:
      • $\bullet$
        What threat models or risk scenarios is the evaluation addressing?
      • $\bullet$
        What specific harms or misuses is the evaluation attempting to identify?
      • $\bullet$
        What are the specific contexts (e.g. user groups and use cases) in which the AI system will be deployed?
      • $\bullet$
        What are the performance expectations in the intended contexts?
    • $\square$
      Type of evaluation:
      • $\bullet$
        Is the evaluation assessing capabilities, propensities, or risks?
      • $\bullet$
        What form of evaluation methodology is being used (benchmark, red-teaming, human uplift studies, or user studies, etc.)?
    • $\square$
      Continuous evaluations:
      • $\bullet$
        Will the system be subject to ongoing monitoring after this evaluation?
      • $\bullet$
        At what intervals will follow-up evaluations occur?
    • $\square$
      Obsolescence criteria:
      • $\bullet$
        What changes (system updates, deployment context, discoveries about similar systems, or technological advances) would render this evaluation obsolete?
      • $\bullet$
        How long is this evaluation expected to remain relevant without updates?
  • $\square$
    How is it evaluated
    • $\square$
      Description of evaluation setup:
      • $\bullet$
        What general evaluation methodology was employed?
      • $\bullet$
        What were the key components of the evaluation design?
      • $\bullet$
        What evaluation framework or established protocol was followed (if applicable)?
    • $\square$
      Justification of capability-to-goal translation:
      • $\bullet$
        How does the evaluation approach align with the capabilities or risks being assessed?
      • $\bullet$
        How do the evaluation procedures relate to the system’s intended use contexts?
      • $\bullet$
        What is the rationale for how the chosen methods measure the targeted capabilities?
    • $\square$
      Interpretation of scores:
      • $\bullet$
        How should the evaluation results be interpreted?
      • $\bullet$
        What limitations should be considered when interpreting the findings?
      • $\bullet$
        What do different performance levels indicate about the system?
  • $\square$
    Access and resources
    • $\square$
      Level of access:
      • $\bullet$
        Did the auditor have black-box, gray-box, or white-box access?
      • $\bullet$
        What specific system documentation was provided to auditors, including access to system code, architecture, design documents, training data, or previous evaluation results?
      • $\bullet$
        Were any aspects of the system explicitly restricted from auditor access?
    • $\square$
      Resources available:
      • $\bullet$
        What computational resources were allocated to the evaluation?
      • $\bullet$
        How much time was allocated for the full evaluation cycle?
      • $\bullet$
        Was the evaluation timeline sufficient for comprehensive testing?
      • $\bullet$
        How many auditors were involved and what was their time commitment?
      • $\bullet$
        What specialized tools or software were available to the auditors?
      • $\bullet$
        Were resources sufficient to conduct all planned evaluation activities?
      • $\bullet$
        What resource constraints limited the scope or depth of the evaluation?
  • $\square$
    Integrity
    • $\square$
      Selection process:
      • $\bullet$
        How were auditors identified and recruited for this evaluation?
      • $\bullet$
        What criteria were used to select auditors?
      • $\bullet$
        What measures were taken to ensure diversity of perspectives among auditors?
      • $\bullet$
        Was there an open call or nomination process for auditors?
      • $\bullet$
        Were independent third parties involved in the selection process?
    • $\square$
      Conflicts of interest:
      • $\bullet$
        What relationships exist between auditors and the organization being evaluated?
      • $\bullet$
        Have auditors previously received compensation from the evaluated organization?
      • $\bullet$
        Do auditors have financial interests or competing interests related to the evaluated system?
      • $\bullet$
        What measures were taken to mitigate potential conflicts of interest?
      • $\bullet$
        Were potential conflicts of interest publicly disclosed?
    • $\square$
      Compensation and incentive structures:
      • $\bullet$
        What was the compensation structure for auditors (fixed fee, hourly, etc.)?
      • $\bullet$
        What contractual limitations were placed on auditors (non-disclosure agreements, etc.)?
      • $\bullet$
        Were there incentives for identifying system flaws or vulnerabilities?
      • $\bullet$
        Were auditors employed internally or contracted externally?
      • $\bullet$
        Were there any performance-based incentives that might bias results?
  • $\square$
    Review and communication
    • $\square$
      Review:
      • $\bullet$
        What aspects of the evaluation (e.g. methodology, report, findings) were subjected to review?
      • $\bullet$
        Who reviewed these aspects, and what was their relationship to the auditors and system developers?
      • $\bullet$
        What expertise did the reviewers bring to the process?
      • $\bullet$
        How were reviewer disagreements resolved and what changes resulted from the review process?
    • $\square$
      Feedback channels:
      • $\bullet$
        What specific mechanisms exist for providing feedback on the audit?
      • $\bullet$
        How can new information be submitted for consideration after publication?
      • $\bullet$
        Who should be contacted with questions or concerns about the evaluation?
      • $\bullet$
        Is there a plan for addressing follow-up questions about the results?
    • $\square$
      Executive summary:
      • $\bullet$
        Does the summary clearly explain what was evaluated, how, and what was found?
      • $\bullet$
        Are limitations of the evaluation explicitly stated?
      • $\bullet$
        Is the summary accessible to non-technical audiences?
    • $\square$
      Definitions of key terms:
      • $\bullet$
        Are technical terms and evaluation metrics clearly defined?
      • $\bullet$
        Is terminology consistent with industry standards?
    • $\square$
      Access to results:
      • $\bullet$
        Who will receive which parts of the evaluation findings?
      • $\bullet$
        How and when will results be published or distributed?
      • $\bullet$
        What information, if any, will be redacted from public versions?
Appendix D Selection of governance documents
We chose the relevant governance documents to include in the analysis by considering the following three factors:
  1. 1.
    Major jurisdiction: broadly selected by geopolitical and economic centrality, especially in the AI industry (Stanford HAI staff, 2024; Bianzino et al., 2023).
  2. 2.
    Authority of the institution: within each jurisdiction, we chose government and non-government institutions. METR was the exception because their evaluation reports have been highly influential in the field, and often involve co-authorship from researchers in scaling labs and AISIs.
  3. 3.
    Direct influence on model evaluations. For model developers, these tended to be their Responsible Scaling Policies or Frontier Safety Frameworks (see METR, a). For government bodies, we selected documents that tracked with more scale and influence (i.e. more national/federal than local, or more binding than not). The EU Code of Practice is a unique outlier in not being strictly binding but in effect being as influential as it enables presumption of conformity with the Act (see European Union and EU monitor ).
To enable the governance documents to fit Table 3, we abbreviated the names of the documents. Here, we present their official names.
  • Official regulations
    • Brazil: We analyzed the Federal Senate Bill No. 2338/2023 (approved by Senate in 2024, but pending further steps to become law). (Brazilian Federal Senate, 2023)
    • China: Governance of AI models and systems is detailed across various documents, primarily the: Draft AI Law (with scholars comments, translated May 2024, mur ); Internet Information Service Algorithmic Recommendation Management Provisions (with effect from 2022, Creemers et al. (2022)); Measures for the Management of Generative Artificial Intelligence Services (Draft for Comment) 2023 (Huang et al., 2023)); China 2022 rules for deep synthesis (Gu, 2023)); and China (Draft Measure on Ethical Review).
    • EU: We included the keystone AI Act (EU AI Act, 2024) with its accompanying Code of Practice (“CoP”) (European Commission, 2025): The EU AI Act needs to be read in conjunction with the Code of Practice. The Code of Practice is a unique type of document which, while technically not legally binding, enables presumption of conformity with the EU AI Act. In effect, due to the legal uncertainty of not knowing what alternatives would comply, it is likely to be followed by those wishing to be compliant with the EU AI Act, resulting in effectively a binding effect until “harmonized standards” are published.
    • Singapore: We analyzed the Model AI Governance Framework, which is not technically binding but authoritative guidance. Model AI Governance Framework for Generative AI (Singapore, 2020)
    • South Korea: We analyzed the “Basic Act on the Development of Artificial Intelligence and Establishment of Foundation for Trust”, which comes into force Jan 2026. Governance of AI models and systems appears to be done by a keystone AI Act (with effect from 2026) (National Assembly of the Republic of Korea, 2024).
    • US: Given the lack of a US federal-wide AI regulation, we examined President Biden’s Executive Order (The White House, 2023), which provided the leading signal as to how the US might think about governing model evaluations and risk assessments. It is now repealed, but part of its legacy remains with the highlighted importance of technical assessments and the NIST AI Risk Management Framework.
  • AI company policies
    • Anthropic: Responsible Scaling Policy (Anthropic, 2024)
    • Google DeepMind: Frontier Safety Framework v2 (Google, 2025)
    • Meta: Frontier AI Framework v1.1 (Meta, 2025)
    • Microsoft: Frontier Governance Framework (Microsoft, 2025)
    • OpenAI: Preparedness Framework Beta (OpenAI, 2023)
    • xAI: Draft Risk Management Framework (xAI, 2025)
Appendix E Features underspecified by current EU Code of Practice stipulations about the Model Report
The EU Code of Practice (“CoP”), if ratified, will be the default way to comply with the AI Act [^12] In theory, model providers can showcase compliance through other means that need to be shown to be equivalent to the Code of Practice. However, given legal uncertainty, this presents a less likely avenue for model developers.. As of its Third Draft version, the Code’s Safety and Security section Commitment II.8 details what Model Reports to be submitted to the AI Office are to include. It covers many aspects, including what our paper calls the core principles and specific features (expertise and background of auditors, scope, methodology, and resources involved in model evaluations). However, it does not specify that the Model Report itself must contain information about channels for external review and feedback for the report, or integrity (conflicts of interest) about the process behind the creation of that report.
These exclusions may be understandable due to varying demands and industry concerns of heavy-handedness (Haeck, 2025). Still, it is worth noting as it allows a clear conversation of whether these issues should be codified, where that code should live, and whether more jurisdictions should adopt this approach (Bradford, 2020). These views will have to be updated with regulatory shifts.
Feature: Review and communication While the AI Office is to have access to key pieces of information like the Model Report, unredacted frameworks, models and evaluation results, channels for non-AI Office peer review are not spelled out. There are multiple candidates for who counts as a relevant “peer” (other model auditors, companies, researchers), and issues, including who would be a trusted coordinator of these interactions, different actors’ contexts and timelines for work, preventing test data leakages and evaluation-gaming. Ironing out the difficulties of this implementation is the subject of future work.
Feature: Integrity (Conflicts of interest) It appears to require appropriate checks for conflicts of interest (Safety and Security section glossary definition of “independent external”) when external auditors are involved, but not the explicit reporting of this in the submitted Model Report.
While we have not included this as part of our audit cards, we note other experts calling for good publishing norms for evaluation reports, which includes: evaluations must be reported per risk and in relation to any risk identification and mitigation efforts by the developer, why certain statistical validity measures were chosen, and how meaningful are the differences between newly tested versus older models or human baselines (see interpretation and quantitative results in Reuel et al. (2024b); Salaudeen et al. ). This non-conclusive list are ways to help others understand why an evaluation finding matters, making it harder to take an evaluation out of context, and easier for a wider community to understand existing coverage of risk assessments, plan for future assessments and how to improve them.
Appendix F Stakeholder interviews
F.1 Interviewee details
Participant Organization Role
P1 Third-party evaluator Evaluation designer, developer & writer
P2 [anonymized] [anonymized]
P3 Scaling lab Evaluation developer
P4 Third-party evaluator Evaluation developer
P5 Third-party evaluator Evaluation developer & writer
P6 [anonymized] Evaluation policy researcher
P7 Third-party Evaluation policy researcher
P8 Government Evaluation policymaker & writer
P9 Third-party evaluator Evaluation development operations support
P10 [anonymized] [anonymized]
Table 5: Overview of the stakeholders interviewed for the semi-structured interviews. The organization and role has been reported up to the level of detail interviewees were comfortable with. Note: Participants P2 and P10 were excluded from the final analysis as their interviews were conducted off the record and attributions could not be confirmed in time for publication.
F.2 Interview guide
The following is our interview guide and set of questions we used for the 30-minute semi-structured interview with each interviewee.
Introductions [3 mins]
  • Introduce our project: We defined what audit cards are, what we did and did not include in that scope
    • Audit cards are a way to standardize reporting of AI evaluation process, with the goal of contextualizing what evals can do and transparency about the process
    • We state examples of what we include: why these metrics, who are the auditors, resources available (compute, time) during eval, how they’re selected and potential COI
    • We disambiguate them from examples we don’t include: specific metrics, technical implications, elicitation techniques;
  • Outline process for interview: We walked interviewees through what to expect of the interview
    • Contribution: Used to color in details about evals process from stakeholders.
    • Privacy and permissions: We checked beforehand what level of recording, attribution, and other process checks they would be comfortable with. We confirmed nothing they said to us will be shared to others without it being run past them and them confirming in writing.
General questions to ask everyone [10 mins max]
  1. 1.
    Describe the work you do in relation to evals, try to be specific about your responsibilities, identities of people/orgs you relate to, your role in the evals field
  2. 2.
    Who is the evaluation report you [build/analyze/use] for? Who is the target audience? E.g. developers (internal, external), research, governance, or also the general public.
  3. 3.
    What part of the evaluation process which, if not done well/reported well, would make you doubt the quality of the eval?
  4. 4.
    How do imagine evaluation best practices best becoming reality? Do you see codification—whether in law or Industry norm—as a good/valid way to quality control?
  5. 5.
    In your view, what are the biggest challenge preventing evals from being better? More useful for improving safety, or whatever their key goal is?
Questions to ask specific categories of people [12 mins]
Evaluation developer
  1. 6.
    Do you consider the limitations and assumptions of your evals process? Where does that thinking get captured (if it does)?
  2. 7.
    What is the background/expertise of auditors—does it depend on the type of evaluation or something else, and who makes that decision?
  3. 8.
    Can you give us examples of evals where the evals process differed and whether it affected the quality of eval?
  4. 9.
    Do you follow any internal best practices?
  5. 10.
    What is the current training process both internally (with evaluation developers, human baseliners) and with your engagement partner?
Evaluation report writer
  1. 11.
    What is the most important information to share in the evaluation report? How do you decide that?
    1. (a)
      Can you walk through two situations in which you decided differently (i.e. type of info/detail level to publish and why)?
    2. (b)
      What’s the right balance of transparency in reporting evals, and what risks surround the achieving of that?
  2. 12.
    From our skim of evaluation reports, we found these tend to be underspecified/not specified—why do you think that might be? e.g. assumptions, auditors’ background, integrity and resources of process, peer review, commitments to take action based on evaluation results, and state criteria that make evaluation obsolete
Evaluation development overseer/manager & translate it to policy people
  1. 13.
    Do you consider the limitations and assumptions of your evals? Where does that thinking get captured (if it does)?
    1. (a)
      e.g. specific downstream application context on which the benchmark is contingent;
  2. 14.
    How do you select the right auditors (external / internal)? Walk through a recent eval’s selection?
  3. 15.
    How much autonomy does your organization get in deciding access/resources (e.g. time/compute) for evaluations [of a lab’s models]? What does this depend on?
    1. (a)
      Do you include different stakeholders in the evals process—who are they, and how are they engaged?
  4. 16.
    What is the current training process both internally (with evaluation developers, human baseliners) and with your engagement partner?
  5. 17.
    What is the right balance of flexibility and specificity in regulation? E.g. what makes an auditor ’qualified’ (background/training), conflict of interest, funding/contract structures
  6. 18.
    If you wanted to, how easy would it be for you to be misleading when communicating evaluation results? Why?
Evaluation operations support (decisions getting access to resources that evaluation developers need)
  1. 19.
    What resources do you usually require for an eval, and what influences that? Do you document resource requirements, and where do you share this?
  2. 20.
    What are the conditions of you being able to access them? E.g. Recent Cyber CBRN Agent evals report used blue/red anonymized models at UK AISI
  3. 21.
    Could you describe the different engagement processes, and why some have been easier than others?
Evaluation used for recommendations - policy researcher/think tank, funders, lobbying, advocacy
  1. 22.
    If you have had to use an evaluation report (could be your org or another org.) to make a recommendation, what features of the report have helped you to do so?
  2. 23.
    Do you have examples of/from evaluation reports you found good and/or bad, and why?
  3. 24.
    How much do you trust evaluation results depending on the funding source (e.g. internal different labs OpenAI, external: Apollo, AISI)
AI safety communicators
  1. 25.
    How do you understand evals (as distinct from audits) and what type of evaluation reporting/communication are you aware of?
  2. 26.
    What insights are people most often seeking from evals, and how are evals (not) meeting that need?
  3. 27.
    Re target audience question: Why do you think [previous answer] is the relevant audience? Other relevant audience segments, how does the messaging change? How can evals most strongly communicate their goal?